OpenStudy (rock_mit182):

a family has two cars. the first car has a fuel efficiency of 40 miles per gallon of gas. during one particular week, the two cars went a combined total of 1800 miles, for a total gas consumption of 65 gallons. how many gallons were consumed by each of the two cars that week?

OpenStudy (rock_mit182):

a family has two cars. the first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiencyof 20 miles per gallon of gas. during one particular week, the two cars went a combined total of 1800 miles, for a total gas consumption of 65 gallons. how many gallons were consumed by each of the two cars that week?

OpenStudy (rock_mit182):

X = number of gallons of car 1 y = number of gallons of car 2 1)\[40\frac{ miles }{ gallon }(x* galllon) 20\frac{ miles }{ gallon }(y*gallon)=1800miles\] 2) x + y = 50 gallons

OpenStudy (nurali):

Equation: 1st car DATA: rate = 40 mi/gal ; # of gallons = x gal ; distance = 40x miles 2nd car DATA: 20 mi/gal ; # of gallons = 65-x gal ; distance = 20(65-x) miles Equation: 40x + 20*65-20x = 1800 20x = 1800-1300 x = 25 gallons (gas consumed by 1st car) 65-x = 40 gallons (gas consumed by 2nd car)